OpenStudy (clara1223):

find the limit as x approaches 121 of (sqrt(x)-11)/(x-121). I have already solved and gotten 1/22 which was right but I can't remember how and can't find my notes. All I need is an explanation

OpenStudy (clara1223):

\[\lim_{x \rightarrow 121}\frac{ \sqrt{x}-11 }{ x-121 }\]

OpenStudy (amistre64):

can we factor the bottom?

OpenStudy (clara1223):

if x was squared then yes but it's not.

OpenStudy (amistre64):

oh but it is ... let a^2 = x, then a = sqrt(x)

OpenStudy (clara1223):

wait I don't understand that, can you explain it further?

OpenStudy (amistre64):

also, as a general rule, if the top and bottom have the same zero ... the zeros cancel out due to a common factor.

OpenStudy (amistre64):

at x=11, we have 0/0 right?

OpenStudy (clara1223):

do you mean at x=121?

OpenStudy (amistre64):

its just a substitution ... if the x does not look like its a square ... replace it by something that looks squarey yeah, x = 121 ... brain is faster then my fingers sometimes

OpenStudy (clara1223):

how would you go about replacing it with something that looks "squarey"?

OpenStudy (amistre64):

if we let x=a^2, and take the sqrt of each side ... sqrt(x) = a a-11 --------- a^2 - 121 now it looks squared to me

OpenStudy (clara1223):

Oh! that helps a lot. So then can factor the denominator to (a+11)(a-11) and cancel a-11 from both sides so im left with 1/a+11?

OpenStudy (amistre64):

your notes might say; divide it al by the highest power of x tho

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

and since sqrt(x) = a sqrt(121) = 11 and we get your coveted 1/22

OpenStudy (clara1223):

Thank you!

OpenStudy (amistre64):

factoring seems to be the best route to me ... and youre welcome again :)